OpenStudy (anonymous):

2x + 10y = -42 -3x + 9y = -81 ELIMINATION METHOD SOMEONE HELP

OpenStudy (anonymous):

Multiply the 1st equation by 3 and the other by -2 and solve for y.

OpenStudy (anonymous):

Multiply the first one by 3. Tell us what do you get. Multiply the second one by 2. Tell us what do you get

OpenStudy (anonymous):

multiplied by 3 got 16 multiplied the second by 2 and got 24

OpenStudy (anonymous):

Multiply the first one by 3 3(2x+10y)=-42 =6x+30y=-126.....(A) and Multiply the second one by -2 -2(-3x+9y)=-81 =6x-18y=-162......(B) then use both eqs (A) n (B) and eliminate the common i.e 6x 6x+30y=-126 6x-18y=-162 - + =+ (sign will be changed) 6x will be cancel out and the eq will become as follow 48y=-288 y=-288/48 y=-6 hope u understand it

OpenStudy (anonymous):

and by placing the value of " y" in one the equation you will get the value of " x" 2x+10y=-42 2x+10(-6)=-42 2x-60=-42 2x=60-42 2x=18 x=18/2 x=9